OpenStudy (anonymous):

Create your own unique literal equation and solve for one of the variables. Show your work.

OpenStudy (saifoo.khan):

three times x equals nine.

OpenStudy (saifoo.khan):

3x = 9 x = 9/3 x = 3

OpenStudy (saifoo.khan):

what do u think?

OpenStudy (anonymous):

haha would that work?

OpenStudy (saifoo.khan):

what does that mean by LIteral Equation?

OpenStudy (anonymous):

4 sailors are marooned on an uninhabited island. They're all out of food. They decide the heaviest sailor will be slaughtered and eaten. He weighs 240 lbs. How many pounds of meat do the other sailors get to eat? 3x = 240 x = 240/3 x=80 lbs

OpenStudy (anonymous):

I guess literal means words? ;)

OpenStudy (anonymous):

@HerbvanH I am your biggest fan!!! That was hysterical! @Saifoo V = π r2h for h is an example

OpenStudy (saifoo.khan):

V = π r2h